Circular No. 6659

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GS 1843+00 A. Santangelo, S. Giarrusso, S. Del Sordo, G. Cusumano, S. Piraino,

and A. Segreto, Istituto di Fisica Cosmica e Applicazioni Informatica, Palermo; D. Dal Fiume and M. Orlandini, Istituto Tecnologie e Studio Radiazioni Extraterrestri, Bologna; A. N. Parmar and T. Oosterbroek, Space Science Department, European Space Agency, ESTEC; and L. Piro and M. Feroci, Istituto di Astrofisica Spaziale, Frascati, report: “A BeppoSAX Narrow Field Instrument observation of the transient x-ray pulsar GS 1843+00 was carried out on Apr 4.0798–4.6250 UT. MECS and LECS observed the source at α = 18^{h}45^{m}34^{s}, δ = +0^{o}52′.5 (equinox 2000.0). The estimated error, which is mainly due to systematic uncertainty, is 30′′. A search for the unidentified optical counterpart is urged. Pulsations of ∼ 29.5 s were clearly detected from the source. The source fluxes measured were as follows (in units of 10^{−10} erg cm^{−2} s^{−1}): 2–10 keV (LECS and MECS), 4.44; 4–40 keV (HPGSPC), 22.1; 20–200 keV (PDS), 15.3. The energy spectrum is well fitted at higher energies (HPGSPC/PDS) by a cutoff power law with photon index 0.39 ± 0.09 and cut-off energy 20.2 ± 1 keV, while at low energies the spectrum shows strong photoelectric absorption with n_{H} = (2.61 ± 0.13) × 10^{22} cm^{−2}. Inspection of the XTE ASM lightcurves reveals that the source was still active at a 2–10-keV flux of 12 mCrab as of May 1.”

COMET C/1997 J2 (MEUNIER-DUPOUY) Total magnitude estimates (visual unless otherwise noted): May 7.99

UT, 13.3 (G. Comello, Roden, The Netherlands, 0.30-m reflector; initially thought to be C/1997 J1); 9.91, 12.6 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 10.03, 12.8 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 10.89, 13.3 (H. Mikuž, Crni Vrh Observatory, Slovenia, 0.20-m f/2 Baker-Schmidt camera + V filter + CCD; tail ∼ 2′ long in p.a. 350^{o}); 10.94, 12.7 (Meyer).

COMET C/1997 J1 (MUELLER) Total magnitude estimates (visual unless otherwise noted): May 7.92

UT, 12.9 (Meyer); 9.95, 13.0 (Bouma); 10.85, 13.5 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 10.87, 12.6 (Mikuž, CCD + V filter); 10.94, 12.2 (Meyer).

1997 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green